OpenStudy (anonymous):

Simplify. Square root a^2/b^5 plz help

OpenStudy (anonymous):

if just the top, then it's a/b^5

OpenStudy (anonymous):

sqrt cancels out the squared.

OpenStudy (anonymous):

square*

Parth (parthkohli):

\( \color{Black}{\Rightarrow \Large \sqrt{a^2 \over b^5} = {(a^2)^{1 \over 2} \over (b^5)^{1 \over 2}} = {a \over b^{2.5}} }\)

Parth (parthkohli):

If it's just the top - \( \color{Black}{\Rightarrow \Large {(a^2)^{1 \over 2} \over b^5} }\) \( \color{Black}{\Rightarrow \Large {a \over b^5} }\)

OpenStudy (anonymous):

\[\frac{\sqrt{a^2}}{\sqrt{b^5}}*\frac{\sqrt{b}}{\sqrt{b}}\] \[\frac{a \sqrt{b}}{b^3}\]

OpenStudy (anonymous):

It's like the first example

OpenStudy (anonymous):

Okay :-).